OpenStudy (anonymous):

Calculate the discriminant. x^2 - 2x + 2 = 0

OpenStudy (accessdenied):

Do you know the formula for the discriminant?

OpenStudy (anonymous):

nope.

OpenStudy (accessdenied):

\( D = b^2 - 4ac \) Using the coefficients from the quadratic form: \(ax^2 + bx + c = 0\)

OpenStudy (anonymous):

so...whats the answer?

OpenStudy (anonymous):

nvm got it.

OpenStudy (accessdenied):

Okay. I get: \((-2)^2 - 4(1)(2) = 4 - 8\\ = -4\)

OpenStudy (anonymous):

thank you :)

OpenStudy (accessdenied):

You're welcome. :) In case you didn't know, if you know the quadratic formula: \[ \frac{-b \pm \sqrt{\color{green}{b^2 - 4ac}}}{2a} \] The green part is the discriminant. It's useful to evaluate the discriminant since it can give us insight on the quadratic's solutions. Just some extra information. :P